### v4

@@ 1 Introduction
Stereo matching represents a fundamental challenge in the field of computer vision, with a range of practical applications including autonomous driving, remote sensing, robotics and augmented reality [1, 2, 3]. The goal of stereo matching is to compute the pixel-by-pixel correspondences between two images, and to generate a disparity map that can be translated into depth using the settings of the stereo camera system [4]. In practical, it needs to achieve excellent matching capabilities in a variety of scenarios.
@@
Since the development of deep learning, numerous stereo matching Convolutional Neural Networks (CNNs) have been presented. Recently, recurrent-based methods have shown promising performance on stereo matching benchmarks. RAFT-Stereo [5] uses a Recurrent All-pairs Field Transform (RAFT) [6] paradigm based on multi-level convolutional Gated Recurrent Units (convGRU). IGEV-Stereo [7] proposes the Combined Geometry Encoding Volume (CGEV) instead of the All-pairs Correlation (APC) Volume [5] in RAFT-Stereo. MoCha-Stereo [8] mining the recurrent feature in the feature channels and achieves state-of-the-art (SOTA) performance. While the limited receptive field restricts the learning of long-range information.
The development of Vision Transformer (ViT) [9] has addressed the issue of acquiring long-range dependencies. STTR [10] uses a transformer architecture, replacing cost volume construction with dense pixel matching from a sequence-to-sequence correspondence perspective. ELFNet [11] combines stereo transformers with a multi-scale cost volume. However, existing stereo transformers still have certain limitations.
@@ 2.1 Recurrent-based Stereo CNNs
Learning-based methods [13, 14, 8] have grown rapidly in recent years. Among them, recurrent-based algorithms stand out for their performance in real-world scenarios. As a representative of recurrent-based techniques, RAFT-Stereo [5] validates the effectiveness of the RAFT iteration [6] and the APC pyramid structure. Subsequent improvements mainly focus on the computation of matching costs in the stereo matching field. CREStereo [15] designs Adaptive Group Correlation Layer (AGCL) to compute the correlation of each feature map at different cascade levels separately, refining the differences independently through multiple iterations. IGEV-Stereo [7] computes an additional set of Geometry Encoding Volumes (GEV) for correlation caculation. MoCha-Stereo [8] restores geometric details lost in the feature channels and calculates a correlation based on recurrent patterns within the channels. This process results in a new cost volume termed Motif Channel Correlation Volume (MCCV). The utilization of MCCV has helped recurrent-based stereo CNNs achieve SOTA performance on many datasets.
Despite the leading performance of such methods, recurrent-based schemes typically use convolutional operations with relatively small receptive fields. They are not conducive to learning global information [9, 16]. This makes recurrent-based CNNs challenging in areas where the semantic interpretation cannot be determined from local geometric information alone.
@@ 2.2 Stereo Transformers
Benefiting from the global learning capability of ViT [9], stereo transformers [10, 11, 17] deliver excellent performance. STTR [10] constructs a transformer module with alternating self-attention and cross-attention to compute feature descriptors. ELFNet [11] utilizes self-attention transformer to learn global features as part of estimating evidence-based disparities. CroCo-Stereo [17] stacks SA and MLP-designed encoder and decoder. Since both approaches incorporate vanilla self-attention (SA) within their transformer blocks, existing stereo transformers struggle to match ill-posed regions such as reflective surfaces and thin structures.
Though the design of the transformer allows for long-range matching [18, 9, 19], it is a computationally expensive model. SA has quadratic time complexity. As a result, existing stereo transformers generally require longer inference times. Currently, there are efforts to reduce the time complexity of stereo transformers. DLNR [20] and RetinaStereo [21] borrow ideas from the attention in Restormer [22, 23] and the recurrent update in RAFT-Stereo [5]. It deviates from the vanilla spatial feature matmul product and instead designs a channel feature matmul product. However, like other recurrent-based methods, DLNR and RetinaStereo lack the spatial interactions of vanilla SA. In addition, although the number of channels $c$, is much smaller than $h\times w$, $c^{2}$ is still much larger than $h\times w$ in DLNR and RetinaStereo. This means that the SA computation of DLNR and RetinaStereo [20, 21] still have a time complexity of $O(c^{2})$.
@@ 3.2.1 Hadamard Product between Query and Key
Vanilla SA [9] in the stereo transformer [10, 11, 17] relies on matmul product to compute the attention matrix $A$. However, this computational approach has a time complexity of $O(n^{2})$. $n=h\times w$ here, $h$ denotes the height, $w$ means width of the feature map. $Q,K,V,A$ stand for query, key, value and attention matrix respectively. In existing stereo transformers, it holds that $n>>c$, where $c$ means the number of feature channels. The above conclusion is a proof that the vanilla SA exhibits a quadratic level of complexity. Despite the introduction of channel-wise SA [22, 20] for stereo transformers, the condition $c^{2}>>n$ still prevails. To improve the computational efficiency of attention, we compute the attention matrix using the Hadamard product, as illustrated in Equ. 2.
@@ 3.2.2 Dense Attention Kernel (DAK)
$V_{i+1}=MKOI(DAK(A_{i})\odot V_{i})\triangleq$
$MKOI(V_{i}+ELU(A_{i})\odot V_{i})$ (5)
@@
$\frac{\partial loss}{\partial V_{i}}=\frac{\partial loss}{\partial V_{I}}\cdot\frac{\partial V_{I}}{\partial V_{i}}\triangleq$
$\frac{\partial loss}{\partial V_{I}}\cdot MKOI(1+\frac{\partial}{\partial V_{i}}\sum_{j=i}^{I-1}(ELU(A_{j})\cdot V_{j}))$ (6)
@@ 3.3 Multi Kernel & Order Interaction (MKOI)
Vanilla self-attention (SA) [9] satisfied $SA(Q,K,V)=SoftMax(\frac{QK^{T}}{\sqrt{d_{k}}})V$. This process involves spatial and channel interactions. Vanilla Hadamard product cannot achieve this. To improve the spatial and channel interaction of HPSA, we propose the Multi Kernel & Order Interaction (MKOI) module as a complement to HPSA.
@@
$\underline{\odot w^{c\rightarrow\frac{c}{2^{m}}}_{s\times s}(V)}]_{m=\{0,1,2\}})$ (7)
$V_{i+1}=HPSA(Q_{i},K_{i},V_{i})$
@@ 3.4 Simple Gate Feed Forward (SGFF)
To complete the transformer architecture and further improve nonlinear expressiveness, we introduce a forward propagation architecture with reference to [22]. This process can be written as Equ.8.
$Gate(V_{i+1})=GELU(w^{c\rightarrow c}_{3\times 3}(V_{i+1}^{in}))\odot w^{c\rightarrow c}_{3\times 3}(V_{i+1}^{in})$
@@
where $GELU$ means GELU operation. The overall computation process of our transformer block can be represented as Equ. 9.
$Block_{att}(V^{init})=V^{init}+HPSA(Q,K,V)$
@@ 3.5 IGEV Decoder
$\mathbb{C}_{GEV}(d,h,w,g)=\mathbb{R}(\frac{1}{N_{c}/N_{g}}\langle$
$f^{g}_{l,4}(h,w),f^{g}_{r,4}(h,w+d)\rangle,F(l))$ (10)
@@
Table 3: Results on the KITTI 2012 Reflective leaderboard. “Noc" means percentage of erroneous pixels in non-occluded areas, “All" means percentage of erroneous pixels in total.
@@
Table 4: Results on the KITTI 2015 and 2012. Error threshold is 3 px for 2015, and 4 px for 2012. Front-ground error is indicated by "F.G.", and Background error by "B.G.". $\star$ refers to Stereo Transformer methods, which are summarized in the lower half of the table. The upper half of the table describes Stereo CNN methods. “Time" denotes the inference time on single NVIDIA Tesla A100.
@@
4 Experiment
4.1 Implementation Details
4.2 Comparisons with State-of-the-art
@@ F H F H F H F H
Table 8: Ablation study. Error threshold is 3 px. “ID" means IGEV Decoder. The following metrics represent the results obtained from training and testing on the Scene Flow dataset. $\frac{Rank(A)}{m}$ computes the average percentage of the ratio between the rank of the attention matrix and its dimension $m$ (the number of rows in the attention matrix) across inference steps. “Time" denotes the inference time on single NVIDIA Tesla A100.
@@
Table 9: Ablation studies on the MKOI module. Experiment No. 1 corresponds to the default configuration of MKOI. The table under "Value Matrix" summarizes the convolution operations along with their corresponding input and output channel dimensions.
@@
4.3 Zero-shot Generalization
4.4 Ablations
Despite achieving results surpassing the CNN paradigm using the vanilla attention mechanism [9], the square complexity of matrix multiplication significantly slowed down the inference time, rendering it unsuitable for efficient matching requirements in practical scenarios. To address this issue, we propose an improved version of transformer encoder called HPSA, and utilize a series of novel designs to ensure its performance.
@@ 4.4 Ablations
Multi Kernel & Order Interaction (MKOI). The HPSA without MKOI can be represented by Equ. 11.
[…]
Since the interaction between matrices is performed entirely through Hadamard product, Equ. 11 fails to facilitate spatial and channel dimensional interactions. We further developed MKOI to overcome this shortcoming. Experiments No.5 and No.6 demonstrate the performance of MKOI.
@@
Table 11: Zero-shot performance of HART-DLNR. The baseline here is DLNR. Our proposed HPSA Is used to replace self-attention mechanism (CATE) of DLNR. All models are trained on Scene Flow, and tested on Middlebury and KITTI. The threshold is 2px for Middlebury, and 3px for KITTI. “Time (s)" is the inference time required for the model to run on the KITTI dataset using the NVIDIA A100.

### v5

@@ Abstract
Hadamard Attention , Stereo Matching , Linear Transformer
journal: Elsevier
@@
[…]
[npu]organization=School of Software, addressline=Northwestern Polytechnical University, city=Xi’an, postcode=710129, country=China
[…]
[pku]organization=Key Laboratory of Integrated Microsystems, Shenzhen Graduate School, addressline=Peking University, city=Shenzhen, postcode=518055, country=China
[…]
[scut]organization=School of Computer Science and Engineering, addressline=South China University of Technology, city=Guangzhou, postcode=510641, country=China
@@ 1 Introduction
Stereo matching represents a fundamental challenge in the field of computer vision, with a range of practical applications including autonomous driving, remote sensing, robotics and augmented reality [1, 2, 3]. The goal of stereo matching is to compute the pixel-by-pixel correspondences between two images, and to generate a disparity map that can be translated into depth using the settings of the stereo camera system. In practical, it needs to achieve excellent matching capabilities in a variety of scenarios.
@@
Since the development of deep learning, numerous stereo matching Convolutional Neural Networks (CNNs) have been presented. Recently, recurrent-based methods have shown promising performance on stereo matching benchmarks. RAFT-Stereo [4] uses a Recurrent All-pairs Field Transform (RAFT) paradigm based on multi-level convolutional Gated Recurrent Units (convGRU). IGEV-Stereo [5] proposes the Combined Geometry Encoding Volume (CGEV) instead of the All-pairs Correlation (APC) Volume [4] in RAFT-Stereo. MoCha-Stereo [6] mining the recurrent feature in the feature channels and achieves state-of-the-art (SOTA) performance. While the limited receptive field restricts the learning of long-range information.
The development of Vision Transformer (ViT) has addressed the issue of acquiring long-range dependencies. STTR [7] uses a transformer architecture, replacing cost volume construction with dense pixel matching from a sequence-to-sequence correspondence perspective. ELFNet [8] combines stereo transformers with a multi-scale cost volume. However, existing stereo transformers still have certain limitations.
@@ 2.1 Recurrent-based Stereo CNNs
Learning-based methods [10, 6, 11] have grown rapidly in recent years. Among them, recurrent-based algorithms stand out for their performance in real-world scenarios. As a representative of recurrent-based techniques, RAFT-Stereo [4] validates the effectiveness of the RAFT iteration and the APC pyramid structure. Subsequent improvements mainly focus on the computation of matching costs in the stereo matching field. CREStereo [12] designs Adaptive Group Correlation Layer (AGCL) to compute the correlation of each feature map at different cascade levels separately, refining the differences independently through multiple iterations. IGEV-Stereo [5] computes an additional set of Geometry Encoding Volumes (GEV) for correlation caculation. MoCha-Stereo [6] restores geometric details lost in the feature channels and calculates a correlation based on recurrent patterns within the channels. This process results in a new cost volume termed Motif Channel Correlation Volume (MCCV). The utilization of MCCV has helped recurrent-based stereo CNNs achieve SOTA performance on many datasets.
Despite the leading performance of such methods, recurrent-based schemes typically use convolutional operations with relatively small receptive fields. They are not conducive to learning global information [13]. This makes recurrent-based CNNs challenging in areas where the semantic interpretation cannot be determined from local geometric information alone.
@@ 2.2 Stereo Transformers
Benefiting from the global learning capability of ViT, stereo transformers [7, 8, 14] deliver excellent performance. STTR [7] constructs a transformer module with alternating self-attention and cross-attention to compute feature descriptors. ELFNet [8] utilizes self-attention transformer to learn global features as part of estimating evidence-based disparities. CroCo-Stereo [14] stacks SA and MLP-designed encoder and decoder. Since both approaches incorporate vanilla self-attention (SA) within their transformer blocks, existing stereo transformers struggle to match ill-posed regions such as reflective surfaces and thin structures.
Though the design of the transformer allows for long-range matching [15], it is a computationally expensive model. SA has quadratic time complexity. As a result, existing stereo transformers generally require longer inference times. Currently, there are efforts to reduce the time complexity of stereo transformers. DLNR [16] and RetinaStereo [17] borrow ideas from the attention in Restormer [18, 19] and the recurrent update in RAFT-Stereo [4]. It deviates from the vanilla spatial feature matmul product and instead designs a channel feature matmul product. However, like other recurrent-based methods, DLNR and RetinaStereo lack the spatial interactions of vanilla SA. In addition, although the number of channels $c$, is much smaller than $h\times w$, $c^{2}$ is still much larger than $h\times w$ in DLNR and RetinaStereo. This means that the SA computation of DLNR and RetinaStereo [16, 17] still have a time complexity of $O(c^{2})$.
@@ 3.2.1 Hadamard Product between Query and Key
Vanilla SA in the stereo transformer [7, 8, 14] relies on matmul product to compute the attention matrix $A$. However, this computational approach has a time complexity of $O(n^{2})$. $n=h\times w$ here, $h$ denotes the height, $w$ means width of the feature map. $Q,K,V,A$ stand for query, key, value and attention matrix respectively. In existing stereo transformers, it holds that $n>>c$, where $c$ means the number of feature channels. The above conclusion is a proof that the vanilla SA exhibits a quadratic level of complexity. Despite the introduction of channel-wise SA [18, 16] for stereo transformers, the condition $c^{2}>>n$ still prevails. To improve the computational efficiency of attention, we compute the attention matrix using the Hadamard product, as illustrated in Equ. 2.
@@ 3.2.2 Dense Attention Kernel (DAK)
$V_{i+1}=MKOI(DAK(A_{i})\odot V_{i})\triangleq MKOI(V_{i}+ELU(A_{i})\odot V_{i})$ (5)
@@
$\frac{\partial loss}{\partial V_{i}}=\frac{\partial loss}{\partial V_{I}}\cdot\frac{\partial V_{I}}{\partial V_{i}}\triangleq\frac{\partial loss}{\partial V_{I}}\cdot MKOI(1+\frac{\partial}{\partial V_{i}}\sum_{j=i}^{I-1}(ELU(A_{j})\cdot V_{j}))$ (6)
@@ 3.3 Multi Kernel & Order Interaction (MKOI)
Vanilla self-attention (SA) satisfied $SA(Q,K,V)=SoftMax(\frac{QK^{T}}{\sqrt{d_{k}}})V$. This process involves spatial and channel interactions. Vanilla Hadamard product cannot achieve this. To improve the spatial and channel interaction of HPSA, we propose the Multi Kernel & Order Interaction (MKOI) module as a complement to HPSA.
@@
$V_{i+1}=HPSA(Q_{i},K_{i},V_{i})$ (8)
@@ 3.4 Simple Gate Feed Forward (SGFF)
To complete the transformer architecture and further improve nonlinear expressiveness, we introduce a forward propagation architecture with reference to [18]. This process can be written as Equ.9 and Equ. 10.
$Gate(V_{i+1})=GELU(w^{c\rightarrow c}_{3\times 3}(V_{i+1}^{in}))\odot w^{c\rightarrow c}_{3\times 3}(V_{i+1}^{in})$ (9)
@@
where $GELU$ means GELU operation. The overall computation process of our transformer block can be represented as Equ. 11 and Equ. 12..
$Block_{att}(V^{init})=V^{init}+HPSA(Q,K,V)$ (11)
@@ 3.5 IGEV Decoder
$\mathbb{C}_{GEV}(d,h,w,g)=\mathbb{R}(\frac{1}{N_{c}/N_{g}}\langle f^{g}_{l,4}(h,w),f^{g}_{r,4}(h,w+d)\rangle,F(l))$ (13)
@@
4 Experiment
4.1 Implementation Details
4.2 Comparisons with State-of-the-art
@@
Table 3: Results on the KITTI 2012 Reflective leaderboard. “Noc” means percentage of erroneous pixels in non-occluded areas, “All” means percentage of erroneous pixels in total.
@@
Table 4: Results on the KITTI 2015 and 2012. Error threshold is 3 px for 2015, and 4 px for 2012. Front-ground error is indicated by ”F.G.”, and Background error by ”B.G.”. $\star$ refers to Stereo Transformer methods, which are summarized in the lower half of the table. The upper half of the table describes Stereo CNN methods. “Time” denotes the inference time on single NVIDIA Tesla A100.
@@ F H F H F H F H
4.4 Ablations
Table 8: Ablation study. Error threshold is 3 px. “ID” means IGEV Decoder. The following metrics represent the results obtained from training and testing on the Scene Flow dataset. $\frac{Rank(A)}{m}$ computes the average percentage of the ratio between the rank of the attention matrix and its dimension $m$ (the number of rows in the attention matrix) across inference steps. “Time” denotes the inference time on single NVIDIA Tesla A100.
@@ 4.4 Ablations
Despite achieving results surpassing the CNN paradigm using the vanilla attention mechanism, the square complexity of matrix multiplication significantly slowed down the inference time, rendering it unsuitable for efficient matching requirements in practical scenarios. To address this issue, we propose an improved version of transformer encoder called HPSA, and utilize a series of novel designs to ensure its performance.
Table 9: Ablation studies on the MKOI module. Experiment No. 1 corresponds to the default configuration of MKOI. The table under ”Value Matrix” summarizes the convolution operations along with their corresponding input and output channel dimensions.
@@
Multi Kernel & Order Interaction (MKOI). The HPSA without MKOI can be represented by Equ. 14.
[…]
Since the interaction between matrices is performed entirely through Hadamard product, Equ. 14 fails to facilitate spatial and channel dimensional interactions. We further developed MKOI to overcome this shortcoming. Experiments No.5 and No.6 demonstrate the performance of MKOI.
@@
Table 11: Zero-shot performance of HART-DLNR. The baseline here is DLNR. Our proposed HPSA Is used to replace self-attention mechanism (CATE) of DLNR. All models are trained on Scene Flow, and tested on Middlebury and KITTI. The threshold is 2px for Middlebury, and 3px for KITTI. “Time (s)” is the inference time required for the model to run on the KITTI dataset using the NVIDIA A100.